# Data-Driven Logistic Regression Ensembles With Applications in Genomics

### v5

@@ 4 Algorithm
where the weights $u_{j,g}$ in the $L_{1}$ penalty term are given by $u_{j,g}=\alpha\lambda_{s}+\lambda_{d}/2\sum_{h\neq g}|\beta_{j}^{h}|.$ For each model the problem thus reduces to a weighted elastic net optimization where the weights in the lasso penalty depend on the value of the coefficients in the other models. We exploit the multi-convex structure of the objective function to develop a block coordinate descent algorithm (Xu and Yin, 2013).
@@ 5 Simulation Study
We investigate the performance of split logistic regression in an extensive simulation study which examines three scenarios. In Scenario 1 all predictors are equally correlated. In Scenario 2 the correlation between active and inactive predictors is lower than the other correlations. In Scenario 3, the active predictors follow a block-correlation structure.
5.1 Simulation Settings
@@
Scenario 1: Data are generated according to the logistic model
where $\mathbf{x}_{A,i}^{T}$ are the active predictors and $\boldsymbol{\beta}_{A}$ the corresponding regression coefficients. In this scenario all predictors are equally correlated with correlation equal to $\rho=0.2$ , $\rho=0.5$, or $\rho=0.8$, respectively.
Scenario 2: Data are generated from the logistic model of Scenario 1 with correlation $\rho_{1}$ between active and inactive predictors and $\rho_{2}$ for all other correlations, where $\rho_{1}$ and $\rho_{2}$ are taken over all combinations $\rho_{1}\in\{0,0.2,0.5\}$ and $\rho_{2}\in\{0.2,0.5,0.8\}$ such that $\rho_{1}<\rho_{2}$.
Scenario 3: Data are generated from a logistic model with active predictors in $B$ disjoint blocks
where $\mathbf{x}_{b,i}$ are the predictor variables for block $b$ and $\boldsymbol{\beta}_{b}$ the corresponding regression coefficients. Each block contains 25 predictors, so the number of blocks in each setting equals $B=\zeta p/25$. The correlation between predictors in different blocks is given by $\rho_{1}$ while $\rho_{2}$ is the correlation between predictors in the same block. The correlations $\rho_{1}$ and $\rho_{2}$ are taken over all combinations of $\rho_{1}\in\{0.2,0.5\}$ and $\rho_{2}\in\{0.5,0.8\}$ such that $\rho_{1}<\rho_{2}$.
5.2 Methods and Performance Measures
We compare the performance of split logistic regression to the nine methods listed in Section 2. For split logistic regression, we use $G=10$ models, which may be a sub-optimal choice. For RGLM and RF we use the default number of models in their R package implementations, $G=100$ and $G=500$ models respectively.
@@ 5.2 Methods and Performance Measures
$\text{RC}=\frac{\sum_{j=1}^{p}\mathbb{I}(\beta_{j}\neq 0,\hat{\beta}_{j}\neq 0)}{\sum_{j=1}^{p}\mathbb{I}(\beta_{j}\neq 0)},\quad\text{PR}=\frac{\sum_{j=1}^{p}\mathbb{I}(\beta_{j}\neq 0,\hat{\beta}_{j}\neq 0)}{\sum_{j=1}^{p}\mathbb{I}(\hat{\beta}_{j}\neq 0)},$
where $\boldsymbol{\beta}$ and $\boldsymbol{\hat{\beta}}$ are the true and estimated regression coefficients, respectively. Since RGLM and RF with their default number of models tend to use all the predictors, we do not compute their RC and PR. For each configuration, we randomly generate $N=50$ training and test sets and measure average performance of each method over the test sets.
5.3 Results
In Table 5.3 the simulation results are summarized by reporting for each performance metric the average rank of the competitors over all simulation settings. Lower ranks indicate better performance. Detailed simulation results are available in the supplementary material.
From Table 5.3 it can be seen that the split logistic regression methods overall performed best in terms of MR, with the black box ensemble methods RF and RGLM being the closest competitors. RF and the adaptive Lasso had the two best ranks in terms of SP followed by the split regression methods, but the split logistic regression methods outperformed the other methods in terms of SE (particularly for low event probabilities, i.e. $\mathbb{P}(Y=1)=0.2$). The split logistic regression also performed favorably compared to the competitors when the proportion of active variables was high ($\zeta=0.2$ and $\zeta=0.4$). The scenario and correlation level $\rho$ between the predictors did not have a strong effect on the performance of Split-Lasso and Split-EN relative to the competitors. The split logistic regression methods (unanimously) yielded the best RC. This comes at a price in PR, which is approximately on par with the single-model base estimators Lasso and EN.
Table 2: Average performance metrics ranks of the methods over all simulation configurations. The two best results for each criterion are highlighted in bold.
@@ 5.3 Results
=2pt {tabu}lllllll Method MR SE SP TL RC PR
Lasso 5.86 6.04 7.14 5.26 4.70 6.41
EN 4.60 4.20 5.77 3.88 3.00 5.88
Adaptive 8.25 9.69 2.86 7.43 4.50 6.31
Relaxed 7.04 5.42 8.60 8.49 6.11 4.81
MCP 9.29 9.15 9.28 8.60 8.44 4.09
SIS-SCAD 9.94 10.06 8.41 9.56 8.51 2.85
Split-Lasso-10 2.34 2.15 4.30 1.81 2.00 6.08
Split-EN-10 1.95 1.88 3.88 1.45 1.00 7.01
RGLM-100 2.59 3.14 3.33 3.42 -- --
RF-500 3.82 6.10 1.54 5.41 -- --
XGB 10.31 8.17 10.90 10.69 6.74 1.57
@@
The TL ranks in Table 5.3 show that split logistic regression also provided the best fits, whereas RF was outranked by the single-model sparse methods Lasso and EN for this criterion.
[…]
6 Medical Genomics Data Applications
To demonstrate the competitive performance of split logistic regression in practice, we apply split logistic regression and its competitors in Section 2 to ten medical genomics datasets available from the Gene Expression Omnibus (GEO) database. Some characteristics of the datasets are given in Table 6. The first eight datasets involve the classification of different types of cancerous cell tissues, whereas the last two datasets involve the separation of psoriasis and multiple sclerosis cell tissues from adjacent normal cell tissue. All datasets are first pre-processed as described in Section 2 using the choices $p=100,250,500$, and $1,000$, respectively, for the number of retained genes according to their $q$-value.
Table 3: GEO identification (ID) codes, sample sizes ($n$), number of genes and datasets descriptions.
[…]
=2pt {tabu}lrrl GEO ID $\boldsymbol{n}$ Genes Description
GSE20347 34 22,277 Esophageal cancerous cell tissue.
GSE23400 106 22,283 Esophageal cancerous cell tissue.
GSE23400 102 22,477 Esophageal cancerous cell tissue.
GSE5364 29 22,283 Esophageal cancerous cell tissue.
GSE25869 75 27,578 Gastric cancerous cell tissue.
GSE10245 58 54,675 Lung cancerous cell tissue.
GSE5364 30 22,283 Lung cancerous cell tissue.
GSE5364 51 22,283 Thyroid cancerous cell tissue.
GSE21942 27 54,675 Multiple sclerosis cell tissue.
GSE14905 54 54,675 Psoriasis cell tissue
6.1 Classification Performances
[…]
Table 4: Average ranks for MR and TL over the ten gene expression datasets in Table 6 for different number of genes retained after pre-processing. The two best results for each column are highlighted in bold.
[…]
From Table 4 it can be seen that Split-EN and Split-Lasso show the best performance as they achieved the best two average ranks for both MR and TL across all number of genes retained. These methods are thus very accurate classifiers and also yield good predicted probabilities as indicated by their small test-sample losses. The logistic elastic net was the closest competitor for all values of $p$ in terms of MR and TL, while RF and RGLM were the third and fourth best methods in terms of overall rank for both MR and TL. However, RF and RGLM were not competitive with split logistic regression or the logistic elastic net on half of the medical genomics datasets in Table 6. Moreover, RGLM is prone to include irrelevant genes due to the inherently random nature of the method and lack of sparsity.
[…]
@@ 6.1 Classification Performances
=2pt Top 1 Top 3 Low 3 Method MR TL MR TL MR TL Lasso 0 0 0 0 0 1 EN 1 2 5 8 0 0 Adaptive 1 0 1 0 1 1 Relaxed 0 0 0 0 1 8 MCP 0 0 0 0 9 3 SIS-SCAD 0 0 0 0 8 6 Split-Lasso 1 0 6 8 0 0 Split-EN 3 7 9 10 0 0 RGLM 3 0 5 2 1 0 RF 1 1 3 2 0 0 XGB 0 0 0 0 9 9
6.2 Ranking Genes
Beyond the good prediction accuracy of split logistic regression ensembles, the information of the individual models can be exploited to identify gene sets in order of importance. For a split logistic regression solution, let $\mathcal{M}_{g}=\{j:\hat{\beta}_{j}^{g}\neq 0\}$ denote the set of the genes selected for model $g\ (1\leq g\leq G)$. Now, by considering the sets
$\mathcal{A}_{k}=\left\{j:\sum_{g=1}^{G}\mathbb{I}\left(j\in\mathcal{M}_{g}\right)\geq k\right\},\quad 1\leq k\leq G,$
where $\mathcal{A}_{G}\subseteq\mathcal{A}_{G-1}\subseteq\cdots\subseteq\mathcal{A}_{1}$, we can study the distribution of the genes across the individual models in the split logistic regression solution. These sets identify genes related to cancerous cell tissue in order of importance, because genes can only appear in several models if these genes allow to largely reduce the loss function to compensate for the diversity penalty in (17).
As an example, consider again the thyroid cancer dataset (GSE5364) from Section 2. With cross-validated values for $\lambda_{s}$, $\lambda_{d}$ and $G$ on the full dataset, the Split-EN solution for these data consists of $G=10$ sparse and diverse models. For this example, we find that $\lvert\mathcal{A}_{7}\rvert=0,\lvert\mathcal{A}_{6}\rvert=1,\lvert\mathcal{A}_{5}\rvert=5,\lvert\mathcal{A}_{4}\rvert=9,\lvert\mathcal{A}_{3}\rvert=39,\lvert\mathcal{A}_{2}\rvert=88$ and $\lvert\mathcal{A}_{1}\rvert=206$. Unsurprisingly, genes shared by multiple models consistently had the same sign, re-enforcing the understanding of their relationship with the thyroid cancer. To demonstrate that split logistic regression can identify important genes that may be missed by other methods, let us consider $\mathcal{A}_{5}$. This set contains five genes that appear in at least half of the 10 individual models in the Split-EN solution and thus yield an important contribution to the ensemble. Interestingly, four of these genes do not appear in the logistic elastic net model and thus would be considered irrelevant for the separation of the classes according to this method. One of these genes is SLC25A3 which is the only gene that appears in six models and encodes a protein that contains two isoforms, one of which contains an exon that regulates thyroid tissue function (Bhoj et al., 2015). The three other genes are EEF2 , SNHG4 and TCEB2. The genes EEF2 and SNHG4 have been identified as target biomarkers in thyroid cancer (see Cai et al., 2022; Lan, 2022, respectively). Interestingly, TCEB2 has been identified as a target biomarker in kidney cancer (Long et al., 2017), however a deep bidirectional connection has been made between these two types of cancer due to genetic irregularities that are common to both types (Bellini et al., 2022). Finally, none of the genes in $|\mathcal{A}_{5}|$ were in the top 250 genes ranked by RF on the full thyroid cancer data. On the other hand, most of the genes ranked in the top 250 by RF (particularly the ones ranked near the very top) also appear in the Split-EN ensemble.

### v6

@@ 4 Algorithm
where the weights $u_{j,g}$ in the $\ell_{1}$ penalty term are given by $u_{j,g}=\alpha\lambda_{s}+\lambda_{d}/2\sum_{h\neq g}|\beta_{j}^{h}|.$ For each model the problem thus reduces to a weighted elastic net optimization where the weights in the Lasso penalty depend on the value of the coefficients in the other models. We exploit the multi-convex structure of the objective function to develop a block coordinate descent algorithm (Xu and Yin, 2013).
@@ 5 Simulation Study
Following the ADEMP framework Morris et al. (2019), we structure our simulation study to systematically evaluate our proposed method against established alternatives.
5.1 Aims
The primary aims of this simulation study are to:
  • Assess the predictive performance of split logistic regression compared to established sparse and ensemble methods under varying conditions.
  • Evaluate variable selection accuracy of our proposed approach.
  • Investigate how different correlation structures among predictors affect method performance.
  • Examine the impact of sample size, event probability, sparsity level, non-linear effects, and interaction effects on performance outcomes.
5.2 Data-Generating Mechanisms
We investigate five simulation scenarios based on the logistic regression framework as given in Equation (1). Across scenarios, we systematically vary several aspects of the data-generating process: correlation structure among predictors, correlation levels, sparsity level, sample size, class imbalance, and functional form (linear, interaction effects, and non-linear effects).
Base Configuration for All Scenarios:
  • Active coefficient values are randomly generated as $(-1)^{z}u$ where $z\sim\text{Bernoulli}(0.3)$ and $u\sim\text{Uniform}(0,1/2)$
  • Predictors follow multivariate normal distribution with mean zero and unit variance
  • Dimension $p=$ 1,000 with sparsity levels $\zeta\in\{0.1,0.2,0.4\}$ (proportion of active variables)
  • Training sample sizes $n\in\{50,100\}$ with event probability $\mathbb{P}(Y=1)\in\{0.2,0.3,0.4\}$
  • Test sample size $m=$ 5,000 for performance evaluation
  • $N=50$ Monte Carlo replications per configuration
To systematically assess method performance, we consider a range of data-generating mechanisms that span common scenarios encountered in high-dimensional biomedical data. Scenarios 1–3 are designed to explore the effects of different predictor covariance structures—namely, exchangeable correlation (Scenario 1), distinct correlation levels between active and inactive predictors (Scenario 2), and a block-wise (modular) correlation structure often observed in gene expression or genomics data (Scenario 3). To further challenge the methods, Scenario 4 adopts this realistic block-wise structure and incorporates within-block interaction effects, while Scenario 5 similarly uses the block correlation structure but introduces nonlinear (quadratic) effects for a subset of predictors.
Scenario 1: All predictors have equal pairwise correlation. Data are generated according to:
where $\mathbf{x}_{A,i}^{T}$ are the active predictors and $\boldsymbol{\beta}_{A}$ the corresponding regression coefficients. All predictors are correlated with each other, with correlation $\rho\in\{0.2,0.5,0.8\}$.
Scenario 2: Differential correlation between active and inactive predictors. Data follow the same logistic model with correlation $\rho_{1}$ between active and inactive predictors and $\rho_{2}$ for all other correlations, where $\rho_{1}\in\{0,0.2,0.5\}$ and $\rho_{2}\in\{0.2,0.5,0.8\}$ such that $\rho_{1}<\rho_{2}$.
Scenario 3: Block structure among active predictors. Data are generated according to:
where $\mathbf{x}_{b,i}$ are the predictor variables for block $b$ and $\boldsymbol{\beta}_{b}$ are the corresponding regression coefficients. Each block contains 25 predictors, with $B=\zeta p/25$ blocks. Correlation between predictors in different blocks is $\rho_{1}\in\{0.2,0.5\}$, while correlation between predictors in the same block is $\rho_{2}\in\{0.5,0.8\}$, such that $\rho_{1}<\rho_{2}$.
Scenario 4: Block structure with interaction effects. Data are generated from:
$\log\left(\frac{p_{i}}{1-p_{i}}\right)=\beta_{0}+\sum_{b=1}^{B}\mathbf{x}_{b,i}^{T}\,\boldsymbol{\beta}_{b}+\sum_{j,k\in\mathcal{I}}\gamma_{jk}x_{ij}x_{ik},\quad 1\leq i\leq n,$
where $\mathcal{I}$ represents a subset of pairs of active predictors selected for interactions. Specifically, we randomly select $\lfloor p\zeta/10\rfloor$ pairs of active predictors to form interactions. The interaction coefficients $\gamma_{jk}$ are generated using the same mechanism as the main effects: $\gamma_{jk}=(-1)^{z}u$ where $z\sim\text{Bernoulli}(0.3)$ and $u\sim\text{Uniform}(0,1/4)$. The correlation structure remains identical to Scenario 3, with blocks of 25 predictors, between-block correlation $\rho_{1}\in\{0.2,0.5\}$, and within-block correlation $\rho_{2}\in\{0.5,0.8\}$.
Scenario 5: Block structure with non-linear effects. Data are generated from:
$\log\left(\frac{p_{i}}{1-p_{i}}\right)=\beta_{0}+\sum_{b=1}^{B}\mathbf{x}_{b,i}^{T}\,\boldsymbol{\beta}_{b}+\sum_{j\in\mathcal{N}}\delta_{j}x_{ij}^{2},\quad 1\leq i\leq n,$
where $\mathcal{N}$ represents a subset of active predictors selected to have additional non-linear effects. We select $\lfloor p\zeta/5\rfloor$ active predictors to have non-linear components. The non-linear coefficients $\delta_{j}$ are generated as $\delta_{j}=(-1)^{z}u$ where $z\sim\text{Bernoulli}(0.3)$ and $u\sim\text{Uniform}(0,1/4)$. The correlation structure remains identical to Scenario 3, with the same blocking and correlation parameters.
5.3 Estimands
The primary estimands of interest are:
  • The true coefficient vector $\boldsymbol{\beta}$ (for variable selection performance)
  • The true conditional class probabilities $p_{i}=P(Y_{i}=1|X_{i}=x_{i})$ (for prediction performance)
5.4 Methods
We compare the performance of split logistic regression to nine established methods as detailed in Section 3 and Table 1. For split logistic regression, we use $G=10$ models. For RGLM and RF we use their R package default configurations with $G=100$ and $G=500$ models respectively. For random forest variable importance, we implement the unbiased Gini importance measure (Nembrini et al., 2018) to address known biases in the standard Gini importance when predictors are correlated.
5.5 Performance measures
We evaluate methods using the following performance metrics:
Prediction performance metrics:
  • Accuracy (ACC): Proportion of correctly classified observations
  • Area Under the ROC Curve (AUC): Measure of discriminative ability across thresholds
  • Sensitivity (SNS): Proportion of true positive cases correctly classified
  • Specificity (SPC): Proportion of true negative cases correctly classified
  • Test-sample loss (TSL): Average negative log-likelihood (14) on test data
For all classification methods, we use a probability threshold of 0.5 to determine class assignment: observations with predicted probabilities above 0.5 are classified as class 1, and those below as class 0. While this standard threshold enables fair comparison across methods, we note that in practical applications, especially in medical contexts, this threshold could be optimized, e.g. based on the relative costs of false positives versus false negatives. The AUC metric provides a threshold-independent assessment of discriminative performance.
Variable selection performance metrics:
  • Recall (RCL): Proportion of truly active variables identified by the model
    $\text{RCL}=\frac{\sum_{j=1}^{p}\mathbb{I}(\beta_{j}\neq 0,\hat{\beta}_{j}\neq 0)}{\sum_{j=1}^{p}\mathbb{I}(\beta_{j}\neq 0)}$
  • Precision (PRC): Proportion of selected variables that are truly active
    $\text{PRC}=\frac{\sum_{j=1}^{p}\mathbb{I}(\beta_{j}\neq 0,\hat{\beta}_{j}\neq 0)}{\sum_{j=1}^{p}\mathbb{I}(\hat{\beta}_{j}\neq 0)}$
Since RF with its default number of models tends to use all predictors, and XGB builds sequential models on residuals making feature selection interpretation challenging, we do not compute RCL and PRC for these methods.
5.6 Results
Table 5.6 presents the average ranks of the prediction performance metrics across all simulation scenarios. The split-ensemble approaches, particularly Split-EN-10 and Split-Lasso-10, demonstrate remarkable predictive performance across all scenarios. Despite using only 10 base models, these methods consistently outperform ensemble approaches that employ 100 models or more (RE-Lasso-100, RE-EN-100, RGLM-100, and RF-500). Split-EN-10 achieves the best overall performance, achieving the best average rank in ACC, SNS, AUC and TSL metrics across nearly all scenarios, making it particularly valuable for class-imbalanced datasets where both overall prediction accuracy and detection of the minority class cases are critical.
A particularly valuable finding is the consistency of the split-ensemble methods across different data structures. Whether dealing with simple main effects, complex interactions, or non-linear relationships, both Split-EN-10 and Split-Lasso-10 maintain their superior performance. This demonstrates that these methods adapt effectively to various data complexities without requiring scenario-specific adjustments. Importantly, this excellent performance is achieved with just 10 models and without relying on randomization techniques employed by RE-Lasso, RE-EN, RGLM, and RF methods. Instead, split-ensemble methods use a well-formulated objective function that balances model fit with variable selection and diversity in a mathematically sound framework. While RF achieved the highest specificity across all scenarios, this came at a significant cost in terms of sensitivity. In contrast, the split-ensemble methods maintained competitive specificity rankings while excelling in sensitivity measures, offering a more balanced classification approach.
Figure 1 illustrates SNS and SPC performance across 50 random training sets under challenging high-dimensional conditions with class imbalance. The split-ensemble methods demonstrate an impressive balance of both metrics, maintaining high sensitivity while preserving competitive specificity. Despite using only 10 base models, Split-Lasso-10 and Split-EN-10 consistently outperform more complex methods. Notably, the closest competitors, RE-Lasso-100 and RE-EN-100 with 100 base models each, show lower sensitivity distributions with their 75th percentiles roughly corresponding to the split methods’ median performance. While RF and Adaptive Lasso achieve high specificity, they sacrifice sensitivity considerably. The split-ensemble approaches offer the best overall classification balance in this challenging scenario.
Table 5: Average prediction performance ranks of the methods across all simulation configurations for main effects, interactions, and non-linear scenarios. The three best results for each criterion are highlighted in bold.
@@ 5.6 Results
=2pt {tabu}l lllll lllll lllll Main Effects Interactions Non-Linear
Method ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL ACC SNS SPC AUC TSL
Split-Lasso-10 2.68 2.37 5.65 3.11 2.00 2.02 2.09 5.44 2.19 1.94 2.37 2.33 5.75 2.48 2.46
Split-EN-10 2.06 2.14 4.84 2.06 1.60 1.72 2.00 5.26 1.59 1.67 1.81 1.96 5.10 1.43 1.72
Lasso 7.68 8.04 9.07 8.14 6.94 7.39 7.74 9.23 7.96 6.91 7.28 7.55 9.35 8.06 7.06
EN 6.32 5.91 7.53 6.99 5.52 5.93 5.52 7.76 6.89 5.48 5.87 5.54 7.89 7.00 5.52
Adaptive 10.91 12.47 3.56 9.88 9.76 11.52 13.09 3.26 10.52 10.07 11.41 13.20 2.81 11.15 9.98
Relaxed 8.92 7.21 11.14 9.48 11.18 8.57 6.81 11.13 9.04 10.44 8.35 6.57 11.56 9.00 10.74
MCP 12.48 11.76 12.23 12.58 11.24 11.85 11.15 11.78 12.11 11.09 11.98 11.20 11.37 12.31 11.26
SIS-SCAD 13.28 12.81 11.21 12.51 12.06 13.37 12.35 11.76 12.76 12.46 13.65 12.44 11.22 12.54 12.50
RuleFit 11.54 10.74 12.09 12.90 13.48 11.72 10.43 12.13 12.78 13.69 11.85 9.94 12.48 12.81 13.85
RE-Lasso-100 4.28 4.40 4.66 4.77 3.81 4.16 4.30 4.67 4.67 3.87 3.72 3.83 5.22 4.50 3.52
RE-EN-100 3.35 3.46 4.94 3.81 3.41 3.38 3.20 5.63 3.91 3.56 2.89 2.91 6.26 3.44 3.02
RGLM-100 3.67 4.35 3.69 4.15 4.99 3.98 5.13 3.73 4.31 4.61 4.54 6.05 3.20 4.72 4.89
RF-500 7.09 9.67 2.67 3.15 9.04 8.35 11.54 1.26 4.70 8.98 8.50 11.63 1.19 4.43 8.44
XGB 10.76 9.66 11.72 11.47 9.96 11.04 9.65 11.96 11.57 10.22 10.78 9.83 11.59 11.13 10.04
@@
Table 6: Average variable selection performance ranks of the methods across all simulation configurations for main effects, interactions, and non-linear scenarios. The three best results for each criterion are highlighted in bold.
@@
=2pt {tabu}l ll ll ll Main Effects Interactions Non-Linear
Method RCL PRC RCL PRC RCL PRC
Split-Lasso-10 4.19 6.36 4.26 6.48 4.11 5.85
Split-EN-10 2.67 7.23 2.74 7.50 2.74 7.11
Lasso 8.11 6.64 8.11 7.26 8.11 7.72
EN 6.69 6.12 6.93 5.39 6.93 5.61
Adaptive 9.00 7.30 9.16 7.91 9.37 8.28
Relaxed 9.81 4.90 9.73 5.35 9.52 6.06
MCP 11.42 4.20 11.33 3.97 11.40 3.81
SIS-SCAD 11.58 3.12 11.67 1.53 11.60 1.57
RuleFit 6.39 3.30 6.07 3.61 6.07 2.98
RE-Lasso-100 2.48 10.50 2.39 10.72 2.35 10.81
RE-EN-100 1.01 11.23 1.00 11.09 1.00 11.11
RGLM-100 4.65 7.09 4.61 7.19 4.80 7.07
RF-500 -- -- -- -- -- --
XGB -- -- -- -- -- --
Table 5.6 presents the average ranks of the variable selection performance metrics across all simulation scenarios. The results reveal a clear trade-off between RCL and PRC among the methods. While RE-EN-100 and RE-Lasso-100 achieve the highest recall, they do so at a significant expense of precision, ranking near the bottom (10-11th place) for PRC across all scenarios. This suggests that these methods tend toward excessive variable inclusion when using 100 base models. In contrast, Split-EN-10 demonstrates an impressive balance, ranking among the top three methods for recall while maintaining mid-range precision performance, all with just 10 models. SIS-SCAD, MCP, and RuleFit excel in precision but perform poorly in recall, indicating overly conservative variable selection. Split-Lasso-10 shows competitive recall comparable to traditional sparse methods while offering better precision than randomization-based ensemble methods that use 10 times more base models. This indicates that Split-Lasso-10 is an effective approach for both predictive and interpretable modeling across various data complexity scenarios.
Figure 2 further illustrates the variable selection performance trade-offs under challenging high-dimensional conditions. The plot confirms that RE-Lasso-100 and RE-EN-100 achieve high recall but at a severe cost to precision, indicating they include numerous irrelevant variables. In contrast, Split-EN-10 and Split-Lasso-10 demonstrate a more balanced profile, maintaining competitive recall with substantially better precision than the randomization-based ensembles, despite using only 10 models. The split-ensemble methods’ favorable position in this recall-precision trade-off space highlights their ability to balance identifying true signals while limiting false discoveries, making them particularly valuable when both accurate prediction and meaningful variable selection are required.
[…]
6 Medical Genomics Data Applications
To demonstrate the competitive performance of split logistic regression in practice, we apply split logistic regression and its competitors in Section 3 to ten medical genomics datasets available from the Gene Expression Omnibus (GEO) database. Some characteristics of the datasets are given in Table 6.
Table 7: GEO identification (ID) codes, dataset descriptions, sample sizes ($n$), class distributions, pairing status, and the number of genes retained after matching probes to genes.
[…]
=2pt {tabu}llrrrr GEO ID Description $\boldsymbol{n}$ Class Distribution Paired Genes
GSE20347 Esophageal cancerous cell tissue 34 17 cancer, 17 normal Yes 13,515
GSE23400 Esophageal cancerous cell tissue 106 53 cancer, 53 normal Yes 13,515
GSE23400 Esophageal cancerous cell tissue 102 51 cancer, 51 normal Yes 11,271
GSE5364 Esophageal cancerous cell tissue 29 16 cancer, 13 normal Yes 13,515
GSE25869 Gastric cancerous cell tissue 64 32 cancer, 32 normal Yes 14,476
GSE5364 Lung cancerous cell tissue 30 16 cancer, 14 normal Yes 13,515
GSE10245 Lung cancerous cell tissue 58 40 adeno, 18 squamous No 13,515
GSE5364 Thyroid cancerous cell tissue 51 16 cancer, 35 normal Yes 13,515
GSE21942 Multiple sclerosis cell tissue 29 14 sclerosis, 15 normal No 23,520
GSE14905 Psoriasis cell tissue 54 34 psoriasis, 20 normal No 23,520
The first eight datasets involve the classification of different types of cancerous cell tissues, whereas the last two datasets involve the separation of psoriasis and multiple sclerosis cell tissues from adjacent normal cell tissue. Our benchmark includes both paired samples, where tissue from the same patient contributes to both disease and control groups, and unpaired samples from different individuals. While most classification methods are designed for independent samples, including paired samples in our benchmark is valuable since such designs are common in real-world biomedical applications. This mixture enables evaluation of how classifiers perform across varied experimental designs encountered in practice. All datasets are first pre-processed as described in Section 3 using the choices $p=100,250,500$, and $1,000$, respectively, for the number of retained genes according to their $q$-value.
Each dataset is randomly split $N=50$ times into a training set and test set. For the proportion of training data we considered both 35% and 50%, where proportion 35% is only used if the resulting training set contains at least 20 observations. Each method is applied to the training data and evaluated on the test set using both ACC and TSL. For the split logistic regression methods we use cross-validation (CV) to determine the optimal number of models, with candidates $G=5$, $10$, and $25$. For $p\in\{100,250,500,1{,}000\}$, Table 8 shows the average ranks across all datasets and training proportions. The results reveal that Split-EN-CV consistently ranks among the top three methods across all feature dimensions, particularly impressive when $p\geq 500$ where it either matches or surpasses RE-EN-100 in ACC while using only 5–25 interpretable sparse models compared to RE-EN’s 100 models. In terms of overall performance, Split-EN-CV achieves the second-best average rank (2.96) for ACC, behind only RE-EN-100 (2.75) despite using far fewer models. Detailed results are available in the supplementary material. We note that some cancer datasets may contain paired samples from the same patients, potentially introducing correlation not explicitly modeled in our analytical approach.
Table 8: Average ranks for MR and TL over the ten gene expression datasets in Table 6 for different number of genes retained after pre-processing. The three best results for each column are highlighted in bold.
@@ 6 Medical Genomics Data Applications
For each dataset we computed the average ranks of the methods based on the results for each training proportion and $p\in\{100,250,500,1,000\}$. Table 9 summarizes the results by showing for each method the number of times it belonged to the top one, top three and lowest three average ranks among the ten datasets. Split-EN-CV demonstrates remarkable performance, achieving the top rank in ACC for half of all datasets and in TSL for 3 datasets, while Split-Lasso-CV consistently appears in the top three for 7 datasets in ACC and 5 in TSL. Notably, neither split method ever falls into the worst three performing methods for either metric, highlighting their exceptional stability across diverse genomic datasets. While RE-EN-100 shows decent performance by appearing in the top three for several datasets, it achieves the top ACC rank in none of them despite utilizing 100 models compared to only 10 for our Split methods. Similarly, RE-Lasso-100 requires ten times the number of models to occasionally match the performance of Split methods. Traditional sparse methods generally underperform, with MCP ranking among the bottom three for 6 datasets in ACC, while more complex ensemble approaches like RF-500 and XGB show inconsistent performance, frequently appearing among the worst performers.
@@
=2pt Top 1 Top 3 Low 3 Method ACC TSL ACC TSL ACC TSL Split-Lasso-CV 0 0 7 5 0 0 Split-EN-CV 5 3 6 6 0 0 Lasso 0 0 0 0 0 1 EN 1 1 2 2 0 0 Adaptive 0 0 1 0 2 0 Relaxed 0 0 0 0 0 10 MCP 0 0 0 0 6 1 SIS-SCAD 0 0 0 0 4 0 RuleFit 0 0 0 0 6 10 RE-Lasso-100 2 1 4 4 0 0 RE-EN-100 0 4 8 9 0 0 RGLM-100 0 1 0 2 0 0 RF-500 2 0 2 2 5 4 XGB 0 0 0 0 7 4
Figure 3 presents a comparative assessment of predictive performance using scaled test sample loss (TSL), which is essentially the deviance on test data, for the GSE10245 lung cancer dataset (differentiating between adenocarcinoma and squamous cell carcinoma). This dataset presents a particularly challenging case due to its class imbalance (SCC comprising only 18 samples and being rarer), making it more complex than many of the balanced datasets evaluated in Table 9. For each replication, we computed the scaled TSL as:
Scaled $\text{Scaled TSL}=\frac{\min(\text{TSL})}{\text{TSL}}$
where $\min(\text{TSL})$ represents the minimum test sample loss achieved by any method on that specific random training-test split. This scaling results in values between 0 and 1, where higher values represent superior performance. The Split-Lasso and Split-EN methods demonstrate remarkable efficiency, consistently achieving optimal or near-optimal performance despite leveraging only $G=10$ models. Their distributions are tightly concentrated near 1, indicating these methods either achieved the lowest TSL or came remarkably close across replications. In contrast, the ensembles based on randomization exhibited greater variability and often produced less reliable probability estimates, as evidenced by their substantially lower scaled TSL values. Split-logistic regression’s ability to maintain high-quality probability calibration with minimal ensemble complexity represents a significant advantage for implementation in clinical contexts, particularly for imbalanced datasets like this one where accurate probability estimation is crucial.
Figure 3: Comparison of model performance using scaled TSL (minimum TSL divided by each method’s TSL). Higher values indicate better performance, with 1.0 representing best probability calibration.